# UniF2ace: A Unified Fine-grained Face Understanding and Generation Model

### v4

@@ 1 Introduction
Recently, unified multimodal models (UMMs) have emerged as a vibrant research area enabling cross-modality understanding and generation within a single “any-to-any” framework, marking a significant step toward artificial general intelligence (AGI) (Wu et al., 2024a; Shi et al., 2024; Li et al., 2024a; Zhou et al., 2024; Team, 2024; Xie et al., 2024). Given the central role of faces in daily life, applying this unified paradigm to achieve fine-grained face understanding and generation is essential for developing human-centric AGI. The practical applications are vast and critical: accurate face understanding is pivotal for identity verification (Srinivasan et al., 2024; Roshdy et al., 2024) and human-computer interaction (Liu, 2024; Chowdary et al., 2023), while high-fidelity face generation drives progress in creative industries (Melnik et al., 2024), virtual avatars (Yan et al., 2024), and data augmentation for model robustness (Melzi et al., 2023). These demanding real-world needs urge facial research to push the boundaries of multimodal understanding and generative modeling.
@@ 3 METHODOLOGY
We introduce our unified multimodal model, UniF2ace, designed to model both the understanding and generation of fine-grained facial attributes. Our approach is realized from two perspectives: loss function (Sec. 3.1) and network architecture (Sec. 3.2). Regarding the generation strategy, we recognize that the generation of fine-grained facial attributes is significantly more challenging than understanding tasks, as highlighted in prior studies (Du et al., 2017; Zhou et al., 2024; Xie et al., 2024). To address this, we harness the theory of score matching in discrete diffusion (Lou et al., 2024) and propose the dual discrete diffusion (D3Diff) training strategy, ensuring the meticulous synthesis of facial details. For network architecture, existing UMMs typically focus on dense architectures (Zhou et al., 2024; Xie et al., 2024) or solely on achieving token-level MoE (Deng et al., 2025), lacking the selective integration of instance features. To overcome these limitations, we introduce token-level and sequence-level MoE layers. Distinct MoE modules are designed for generation and understanding tasks, selectively integrating information such as facial embeddings to enhance the model’s ability to capture subtle facial attributes.
@@ 4.1 UniF2aceD-1M Dataset for Fine-grained Face understanding and generation
Dataset Face Resolution VQA Availability Image Caption Tokens (Avg.) Face Attributes (Avg. per Caption) LAION-Face (Zheng et al., 2022) Low 50M 16 2.7 FLIP-80M (Li et al., 2024c) 80M 22 4.4 FFHQ-Text (Zhou & Shimada, 2021) High 760 45 12.2 MM-CelebA-HQ (Karras et al., 2018) 30K 26 6.2 CelebV-Text (Yu et al., 2023b) 70K 80 4.3 UniF2aceD-1M (Ours) (1M) 130K 120 17.7
@@ 4.2 Metrics and Other Facial Datasets
We rigorously evaluate UniF2ace’s performance across both generation and understanding tasks on our UniF2aceD-1M test set. To provide a comprehensive assessment and verify the generalizability of our method, we also conduct evaluations on other public benchmarks, including FFHQ-Text (Zhou & Shimada, 2021), MM-CelebA (Xia et al., 2021), and CelebV-Text (Yu et al., 2023b). For generation tasks, we used VQAscore to measure the relevance of generated images to captions, reporting results based on CLIP-FlanT5-11B (VQAscore-CF5) (Lin et al., 2024b) and LLaVA-v1.5-13B (VQAscore-LV) (Liu et al., 2024c) for robust assessment. We also employ Fréchet Inception Distance (FID) to measure similarity to ground truth and VLM-score to evaluate facial realism. For understanding tasks, we follow LLaVA (Liu et al., 2023) and use GPT-4o (Hurst et al., 2024) and DeepSeek-v3 (Liu et al., 2024a) to score responses on a 1-10 scale across two dimensions: detailed captioning (Desc-GPT, Desc-DS), assessing accuracy in capturing face attributes, and VQA (Conv-GPT, Conv-DS), measuring precision in responding to fine-grained queries. To fully validate UniF2ace, we compare it with SOTA models. This includes generative models such as autoregressive LlamaGen (Yu et al., 2023c) and diffusion-based Stable Diffusion 3 (SD3) (Esser et al., 2024), as well as leading unified multimodal models (UMMs) like TokenFlow (Qu et al., 2024) and OmniFlow (Li et al., 2024a). More implementations details are in Appendix C.
@@ 4.3 Face Generation
Generation Performance on UniF2aceD-1M and Public Dataset. On our UniFaceD-1M benchmark (Tab. 1), our 1.8B parameter UniF2ace sets a new state-of-the-art, outperforming all competing UMMs on key generation metrics including FID, VQA-score, and VLM-score. Furthermore, the model also demonstrates robust generalization, consistently achieving leading scores on public cross-facial datasets such as FFHQ-Text (Zhou & Shimada, 2021), MM-CelebA (Xia et al., 2021), and CelebV-Text (Yu et al., 2023a) (Tab. 2). This strong and consistent performance validates the effectiveness of our D3Diff loss and multi-level grouped MoE architecture for high-quality, fine-grained facial image generation across diverse settings.
@@ Appendix A Related Works
Recent works (Ma et al., 2024; Li et al., 2024d; Wu et al., 2024a; Chen et al., 2025; Wang et al., 2024c) in image understanding and generation have primarily focused on unified multimodal models (UMMs). Early approaches (Li et al., 2024b; Wu et al., 2024b) often integrated external decoders of diffusion models (DMs) with text autoregressive models (ARMs). Inspired by next-token prediction tasks, they proposed using a single Transformer (Vaswani et al., 2017) model to unify understanding and generation (Wu et al., 2024c). For instance, Janus-Pro (Chen et al., 2025) decouples the visual encoder into specialized tokenizers for separate handling of understanding and generation tasks. Chameleon (Team, 2024) and Emu3 (Wang et al., 2024c) employ an ARM to simultaneously manage both tasks, highlighting the advantages of autoregressive models in multitask settings. Additionally, Transfusion (Zhou et al., 2024) and Show-o (Xie et al., 2024) combine a text ARM with a visual DM, enabling seamless integration of image understanding and generation. These studies have advanced the fusion of visual and text generation models, enhancing performance on multimodal tasks. However, despite the proliferation of UMMs, their application has largely been limited to generic domain tasks, with limited exploration in fine-grained visual analysis, particularly in the face domain. Unlike previous UMMs that simply combine ARMs and DMs, we pioneer sparse UMMs by introducing both token-level and sequence-level Mixture of Experts (MoEs), significantly improving model performance.

### v5

@@ 1 Introduction
Recently, unified multimodal models (UMMs) have emerged as a vibrant research area enabling cross-modality understanding and generation within a single “any-to-any” framework, marking a significant step toward artificial general intelligence (AGI) (Wu et al., 2024a; Shi et al., 2024; Li et al., 2024a; Zhou et al., 2024a; Team, 2024; Xie et al., 2024; Zhou et al., 2026; 2024b). Given the central role of faces in daily life, applying this unified paradigm to achieve fine-grained face understanding and generation is essential for developing human-centric AGI. The practical applications are vast and critical: accurate face understanding is pivotal for identity verification (Srinivasan et al., 2024; Roshdy et al., 2024) and human-computer interaction (Liu, 2024; Chowdary et al., 2023), while high-fidelity face generation drives progress in creative industries (Melnik et al., 2024), virtual avatars (Yan et al., 2024), and data augmentation for model robustness (Melzi et al., 2023). These demanding real-world needs urge facial research to push the boundaries of multimodal understanding and generative modeling.
@@ 3 METHODOLOGY
We introduce our unified multimodal model, UniF2ace, designed to model both the understanding and generation of fine-grained facial attributes. Our approach is realized from two perspectives: loss function (Sec. 3.1) and network architecture (Sec. 3.2). Regarding the generation strategy, we recognize that the generation of fine-grained facial attributes is significantly more challenging than understanding tasks, as highlighted in prior studies (Du et al., 2017; Zhou et al., 2024a; Xie et al., 2024). To address this, we harness the theory of score matching in discrete diffusion (Lou et al., 2024) and propose the dual discrete diffusion (D3Diff) training strategy, ensuring the meticulous synthesis of facial details. For network architecture, existing UMMs typically focus on dense architectures (Zhou et al., 2024a; Xie et al., 2024) or solely on achieving token-level MoE (Deng et al., 2025), lacking the selective integration of instance features. To overcome these limitations, we introduce token-level and sequence-level MoE layers. Distinct MoE modules are designed for generation and understanding tasks, selectively integrating information such as facial embeddings to enhance the model’s ability to capture subtle facial attributes.
@@ 4.1 UniF2aceD-1M Dataset for Fine-grained Face understanding and generation
Dataset Face Resolution VQA Availability Image Caption Tokens (Avg.) Face Attributes (Avg. per Caption) LAION-Face (Zheng et al., 2022) Low 50M 16 2.7 FLIP-80M (Li et al., 2024c) 80M 22 4.4 FFHQ-Text (Zhou and Shimada, 2021) High 760 45 12.2 MM-CelebA-HQ (Karras et al., 2018) 30K 26 6.2 CelebV-Text (Yu et al., 2023b) 70K 80 4.3 UniF2aceD-1M (Ours) (1M) 130K 120 17.7
@@ 4.2 Metrics and Other Facial Datasets
We rigorously evaluate UniF2ace’s performance across both generation and understanding tasks on our UniF2aceD-1M test set. To provide a comprehensive assessment and verify the generalizability of our method, we also conduct evaluations on other public benchmarks, including FFHQ-Text (Zhou and Shimada, 2021), MM-CelebA (Xia et al., 2021), and CelebV-Text (Yu et al., 2023b). For generation tasks, we used VQAscore to measure the relevance of generated images to captions, reporting results based on CLIP-FlanT5-11B (VQAscore-CF5) (Lin et al., 2024b) and LLaVA-v1.5-13B (VQAscore-LV) (Liu et al., 2024c) for robust assessment. We also employ Fréchet Inception Distance (FID) to measure similarity to ground truth and VLM-score to evaluate facial realism. For understanding tasks, we follow LLaVA (Liu et al., 2023) and use GPT-4o (Hurst et al., 2024) and DeepSeek-v3 (Liu et al., 2024a) to score responses on a 1-10 scale across two dimensions: detailed captioning (Desc-GPT, Desc-DS), assessing accuracy in capturing face attributes, and VQA (Conv-GPT, Conv-DS), measuring precision in responding to fine-grained queries. To fully validate UniF2ace, we compare it with SOTA models. This includes generative models such as autoregressive LlamaGen (Yu et al., 2023c) and diffusion-based Stable Diffusion 3 (SD3) (Esser et al., 2024), as well as leading unified multimodal models (UMMs) like TokenFlow (Qu et al., 2024) and OmniFlow (Li et al., 2024a). More implementations details are in Appendix C.
@@ 4.3 Face Generation
Generation Performance on UniF2aceD-1M and Public Dataset. On our UniFaceD-1M benchmark (Tab. 1), our 1.8B parameter UniF2ace sets a new state-of-the-art, outperforming all competing UMMs on key generation metrics including FID, VQA-score, and VLM-score. Furthermore, the model also demonstrates robust generalization, consistently achieving leading scores on public cross-facial datasets such as FFHQ-Text (Zhou and Shimada, 2021), MM-CelebA (Xia et al., 2021), and CelebV-Text (Yu et al., 2023a) (Tab. 2). This strong and consistent performance validates the effectiveness of our D3Diff loss and multi-level grouped MoE architecture for high-quality, fine-grained facial image generation across diverse settings.
@@ Appendix A Related Works
Recent works (Ma et al., 2024; Li et al., 2024d; Wu et al., 2024a; Chen et al., 2025; Wang et al., 2024c) in image understanding and generation have primarily focused on unified multimodal models (UMMs). Early approaches (Li et al., 2024b; Wu et al., 2024b) often integrated external decoders of diffusion models (DMs) with text autoregressive models (ARMs). Inspired by next-token prediction tasks, they proposed using a single Transformer (Vaswani et al., 2017) model to unify understanding and generation (Wu et al., 2024c). For instance, Janus-Pro (Chen et al., 2025) decouples the visual encoder into specialized tokenizers for separate handling of understanding and generation tasks. Chameleon (Team, 2024) and Emu3 (Wang et al., 2024c) employ an ARM to simultaneously manage both tasks, highlighting the advantages of autoregressive models in multitask settings. Additionally, Transfusion (Zhou et al., 2024a) and Show-o (Xie et al., 2024) combine a text ARM with a visual DM, enabling seamless integration of image understanding and generation. These studies have advanced the fusion of visual and text generation models, enhancing performance on multimodal tasks. However, despite the proliferation of UMMs, their application has largely been limited to generic domain tasks, with limited exploration in fine-grained visual analysis, particularly in the face domain. Unlike previous UMMs that simply combine ARMs and DMs, we pioneer sparse UMMs by introducing both token-level and sequence-level Mixture of Experts (MoEs), significantly improving model performance.